OpenStudy (anonymous):

Radioactive radium has a half-life of approximately 1599 years. What percent of a given amount remains after 760 years? (Round your answer to two decimal places.)

OpenStudy (anonymous):

Try\[x=x_0(\frac{ 1 }{ 2 })^{\frac{ t }{ T_{1/2} }}\]where x is the amount of remaining undecayed material, x_0 is the initial amount of undecayed material, t is the time, and T_1/2 is the half life.

OpenStudy (anonymous):

I got 72.12%

OpenStudy (anonymous):

I get very close to the same answer, but not exactly.

OpenStudy (anonymous):

\[\frac{ x }{ x_0 } = \left( \frac{ 1 }{ 2 } \right)^\frac{ 760 }{ 1599 } = ?\]

OpenStudy (anonymous):

YES THANK YOU

OpenStudy (anonymous):

You're welcome